Solve the Equation by completing the square. Simplify the answers.
2x^2-8x=1

Answer by mananth(16946)

2x^2-8x =1
/2
x^2-4x=1/2
x^2-4x+4 = 1/2 +4
(x-2)^2= 9/2
take the square root
x-2 = +/- 3/sqrt(2)
x=2+ 3/sqrt(2)
OR
x=2- 3/sqrt(2)